OpenStudy (anonymous):

Double-angle Identities: Solve this equasion for all 3 answers (leaving them in terms of pi) 2sin^2(x) - sin(x) = 1 I got it simplified this far: -sin(x)=1-sin^2(x) -sin(x)=cos(2x) Where would I go from here?

OpenStudy (anonymous):

its not right , I dont really know where u got some of the lines from

OpenStudy (anonymous):

if anything you started with an equation with double angles, and you used identities to come up with the above

OpenStudy (anonymous):

that first line is quadratic equation in sin(x)

OpenStudy (anonymous):

you should solve it like a quadratic

OpenStudy (anonymous):

(I've always been really bad at seeing the steps involved with trig, so I'm not surprised my current steps are wrong.)

OpenStudy (anonymous):

I think I understand what you said about solving like a quadratic though, I shall try now.

OpenStudy (anonymous):

put sin(x) = a then you get \[2a^2-a=1\] \[2a^2-a-1=0\] \[(a-1)(2a+1)=0\] \[a=1\] or \[a=-\frac{1}{2}\]

OpenStudy (anonymous):

now go back to trig \[\sin(x)=1\] \[x=\frac{\pi}{2}\]

OpenStudy (anonymous):

\[\sin(x)=-\frac{1}{2}\] \[x = \frac{7\pi}{6}\] or \[x=\frac{11\pi}{6}\]